OpenStudy (anonymous):

I'm supposed to write the slope-intercept form of the line that passes through the point (1,0) and is parallel to x-y=7? Could you please help me work through this problem?

OpenStudy (jdoe0001):

well, if you solve x-y=7 for "y", what does it give you?

OpenStudy (anonymous):

You would subtract X from both sides and get y=x-7?

OpenStudy (jdoe0001):

hmm yeap, so what is the slope of say y = x-7?

OpenStudy (anonymous):

The slope would be 0?